Use the elimination (addition) method to solve the system of equations. If there are no solution or infinite number of solutions, say so.
2x + 4y = 16
4x + 8y = 32

Answer by ikleyn(52754)

.
2x + 4y = 16    (1)
4x + 8y = 32    (2)


Divide equation (1) by 2 (both sides.
Divide equation (2) by 4 (both sides.  You will get an equivalent system

x + 2y = 8      (1')
x + 2y = 8      (2')


These two equations are identical.
So, actually, you have ONLY ONE independent equation in two unknowns.
It has INFINITELY MANY solutions.


What I did for you here, was not Elimination method.


But it is the shortest and clearest way to explain you about this system.

On Elimination method and closely related other methods see the lessons
    - Solution of a linear system of two equations in two unknowns by the Substitution method
    - Solution of a linear system of two equations in two unknowns by the Elimination method
    - Solution of a linear system of two equations in two unknowns using determinant
    - Geometric interpretation of a linear system of two equations in two unknowns
    - Solving word problems using linear systems of two equations in two unknowns
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".